# Quantum Reservoir Computing for Realized Volatility Forecasting

### v1

@@ II.3 Machine Learning Preliminaries
II.3.1 Feedforward Neural Networks
@@
II.3.2 Long Short-Term Memory Neural Networks
@@
where $W_{f}$ is the weight matrix for the forget gate, $\bm{b}_{f}$ is the bias term, and $\bm{f}_{t}$ is a vector whose elements take values between 00 and $1$. The input gate determines which parts of the current input $\bm{x}_{t}$ should be added to the cell state $\bm{c}_{t}$. It has two components, a sigmoid layer to decide which values to update, and a tanh layer to create new candidate values to potentially add to the cell state $\bm{c}_{t}$, which formula are defined as
@@
where $\bm{i}_{t}$ is the input gate output, a vector of values between 00 and $1$, and $\tilde{\bm{c}}_{t}$ is the candidate cell state. The information from the forget gate and the input gate are used to update the cell state $\bm{c}_{t}$, as
@@
In general, the output $\hat{\bm{y}}_{t}$ is a function of $h_{t}$ as $\hat{\bm{y}}_{t}=g(h_{t})$, where $g(\cdot)$ depends on the specific problem. Training is carried out by optimizing a loss function, of the form of Eq. (20), which aims to minimize the difference between $\hat{\bm{y}}_{t}$ and the real output $\bm{y}_{t}$. During the training procedure, the weight matrices $\{W_{f},W_{i},W_{c},W_{o}\}$ and bias vectors $\{\bm{b}_{f},\bm{b}_{i},\bm{b}_{c},\bm{b}_{o}\}$ are all trained and updated iteratively, see Fig. 1(b). The LSTM architecture effectively mitigates challenges such as vanishing and exploding gradients, which are common in standard recurrent neural networks, and ensures the retention of long-term information. These characteristics make LSTM widely used in fields such as natural language processing and time-series prediction.
II.3.3 Classical Reservoir Computing
@@ III Quantum Reservoir Computing for Realized Volatility Forecasting
Figure 2: Quantum reservoir computing schematic. The task is to forecast $\widehat{RV}_{t}$ based on the past information of features $\bm{x}$ at discrete time from $t{-}3$ to $t{-}1$. The Quantum Reservoir consists of two subsystems, the input qubits $\rho_{I}$ for inputting features and the hidden qubits $\rho_{h}$ for hidden state, composed of $n_{1}$ and $n_{2}$ qubits respectively. In the first, the feature variables $\bm{x}_{t-3}$ are encoded into the input qubits via single-qubit gates $R_{Y}(\bm{x}_{t-3})$ applied on $|0\rangle^{\otimes n_{1}}$. Therefore the input density matrix is given as $\rho_{I}(\bm{x}_{t-3})$. The hidden state at first is initialized as $\rho_{h}(t{-}3)=|0\rangle\langle 0|^{\otimes n_{2}}$. Thus the quantum reservoir state is given by $\rho_{I}(\bm{x}_{t-3})\otimes\rho_{h}(t{-}3)$. The state evolves under the action of the Hamiltonian $H$ for time $\tau$, described by $e^{-iH\tau}$. To update the hidden state for the next step, the input $n_{1}$ qubits are discarded such that the new hidden state is given by a partial trace as $\rho_{h}(t{-}2)$. This process continues until all features from $t{-}3$ to $t{-}1$ have been encoded within the quantum reservoir. After processing all features, measurements on Pauli-Z basis of all qubits are performed to read out useful information from the quantum reservoir, yielding the expectation value $\langle Z_{j}\rangle_{\tau}$ for the $j$-th qubit. These expectation values collectively form the readout vector $\bm{m}_{t}^{(1)}=[\langle Z_{1}\rangle_{\tau},\cdots,\langle Z_{n_{1}+n_{2}}\rangle_{\tau}]$. A linear regression model is used to predict $\widehat{RV}_{t}$ based on this readout vector as $\widehat{RV}_{t}=W_{out}\bm{m}_{t}^{(1)}$, where $W_{out}$ is a weight matrix trained using the ridge regression method.
@@
where $X_{i}$ and $Z_{i}$ are the Pauli operators acting on $i$-th qubits (see Table 1 for the definition of Pauli matrices), $J_{ij}$ are exchange couplings between qubits $i$ and $j$ which are randomly sampled from $J_{ij}\in[0,1]$, and $v=1$ is the strength of the magnetic field. After being randomly initialized, the Hamiltonian $H$ is fixed throughout the process, similar to a classical reservoir.
@@
The goal in this paper is to predict $RV_{t}$ by using past $k$ steps features, i.e., the time series $\bm{x}_{t-k},\cdots,\bm{x}_{t-1}\}$. For convenience of notation, we denote $\bm{x}_{t-k}=[x_{t-k,1},x_{t-k,2},\cdots,x_{t-k,n_{1}}]^{T}$, where each element $x_{t-k,i}$ represents the $i$-th economic feature at time step $t{-}k$. For example the input vector $\bm{x}_{t-k}$ may contain three input data at time $t-k$ such as realized volatility $RV_{t-k}$, Dividend Yield ($DP_{t-k}$) ratio and Earning-Price ($EP_{t-k}$) ratio (note that all external features used in this paper are defined in Table. 2). In this situation, $\bm{x}_{t-k}=[RV_{t-k},DP_{t-k},EP_{t-k}]^{T}$ contains three input features. In this work, we limit $k{=}3$, that is, we consider the memory depth for learning to be only three steps. Although generalizing to other memory depths is straightforward and it may appear that considering large memories may yield better results, one must also consider the fact that larger memory depths require deeper quantum circuits. Given the finite coherence times available in near-term quantum devices, this trade-off quickly becomes considerable. As we shall see, our choice of $k{=}3$ already yields a satisfactory predictive accuracy. Note that all features are scaled between $[-\pi,+\pi]$, fulfilling the requirement of the $R_{Y}$ gate. The input vectors $\{\bm{x}_{t-3},\bm{x}_{t-2},\bm{x}_{t-1}\}$ are fed into the reservoir model sequentially. The reservoir performs an iterative loop consisting of three steps:
@@
where $\text{where }\quad R_{Y}(\bm{x}_{t-3})$ $=$ $\bigotimes_{j=1}^{n_{1}}e^{-i\frac{x_{t-3,j}}{2}Y_{j}}.$ (30)
Therefore, the input density matrix is given by $\rho_{I}(\bm{x}_{t-3})=|\psi_{I}(\bm{x}_{t-3})\rangle\langle\psi_{I}(\bm{x}_{t-3})|$. The hidden $n_{2}$ qubits (the total reservoir thus consists of $n_{1}+n_{2}$ qubits) at this stage are initialized as
@@
Thus, the collective input state of the quantum reservoir state is given by $\rho_{I}(\bm{x}_{t-3})\otimes\rho_{h}(t{-}3)$. Then the whole reservoir evolves under the action of the Hamiltonian $H$ for a specific time $\tau$, described by $U=e^{-iH\tau}$. The evolution scrambles the input data $\bm{x}_{t-3}$ across the entire reservoir. Following this, while the original $n_{1}$ qubits are discarded, the quantum state of the hidden $n_{2}$ qubits carry this information forward to the next step. The quantum state of the hidden qubits after discarding the input qubits now becomes
@@
where $\text{where }\quad R_{Y}(\bm{x}_{t-2})$ $=$ $\bigotimes_{j=1}^{n_{1}}e^{-i\frac{x_{t-2,j}}{2}Y_{j}}.$ (34)
@@
where $\text{where }\quad R_{Y}(\bm{x}_{t-1})$ $=$ $\bigotimes_{j=1}^{n_{1}}e^{-i\frac{x_{t-1,j}}{2}Y_{j}}.$ (37)
@@
where $M_{1}=[\cdots,\bm{m}_{t}^{(1)},\cdots,\bm{m}_{T}^{(1)}]$, $I$ is the identity matrix and $\delta{=}10^{-8}$ is a small number which is used to guarantee that the matrix is non-singular. We call this model as Quantum Reservoir 1 (QR1), since here we only use one quantum reservoir.
@@
Furthermore, to obtain richer information from the quantum reservoir, we adopt the ensemble reservoir approach. In this approach, at any given step $t$ apart from the measurement outcomes $\bm{m}^{(1)}_{t}$, one can use another reservoir which is almost identical to the previous setup except that the evolution at the step III is run for a time duration of $\tau/2$, instead of $\tau$, see Fig. 3. The ensemble of these two reservoir setus are combined to make a larger vector $\bm{m}_{t}^{(2)}=[\langle Z_{1}\rangle_{\tau},\ldots,\langle Z_{n_{1}+n_{2}}\rangle_{\tau},\langle Z_{1}\rangle_{\tau/2},\ldots,\langle Z_{n_{1}+n_{2}}\rangle_{\tau/2}]^{T}$, where $\langle Z_{j}\rangle_{\tau/2}$ represents the measurement of the Pauli operator $Z_{j}$ in the second reservoir setup in which the last evolution is run for time $\tau/2$, as schematically shown in Fig. 3. We call this model Quantum Reservoir 2 (QR2) which is expected to be more precise than the QR1 as it uses twice as much resources.
@@ IV.2 Training and Benchmarking
We employ our quantum reservoir models QR1 and QR2 to predict realized volatility. To assess their performance, we conduct a comprehensive benchmarking study against established classical models. In particular, we compare quantum reservoir computing models QR1 and QR2 with the classical linear models AR1 (which is AR with memory of one previous step), AR3 (which is AR with memory of three previous steps), ARMAX, HAR, HARX as well as the non-linear machine learning based methods LSTM, LSTMX, RC and RCX, see the relevant parts of section II for a short introduction to these models. For all models, including classical models and quantum reservoir computing, we employ a rolling-window approach to estimate and train the models, optimize their parameters, and perform one-step-ahead out-of-sample predictions. The initial training window covers the period from February 1950 to June 1997 (approximately 571 months). After training and optimizing these models on the initial window, we generate a forecast for the next out-of-sample month (August 1997). Subsequently, the rolling window advances by one month, using data from March 1950 to August 1997 to predict September 1997. This process continues iteratively, rolling through all 245 out-of-sample observations, spanning from August 1997 to December 2017. During this process, all models are re-estimated at each step to ensure optimal performance for the new rolling window.
@@
Among the classical linear models that we use, AR1, AR3, ARMAX, HAR, HARX are estimated using ordinary least squares or maximum likelihood methods, depending on the structure of the model. In addition, the machine learning based LSTMX model is implemented with two layers of LSTM cells. The hidden state size is set to 60 for the LSTM model and 50 for the LSTMX variant. The model parameters are optimized using the ADAM optimizer, with a learning rate of $0.001$. For the classical reservoir computing models (RC and RCX), the reservoir consists of 50 hidden neurons for RC and 20 hidden neurons for RCX. The leak rate, which controls the speed at which the state of the reservoir evolves, is set at 0.6. The spectral radius, which influences the stability and non-linearity of the reservoir, is fixed at 0.9. The readout matrix $W_{out}$, responsible for mapping the reservoir states to the output, is estimated using ridge regression to mitigate overfitting and improve generalization. All machine learning hyperparameters, including architecture choices and training settings, are selected using time-series-aware cross-validation procedures within the training window to ensure generalization and robustness.
@@
For quantum reservoir computing, taking into account the current limitations of quantum devices, we used a quantum reservoir with 10 qubits (comprising both input and hidden qubits such that $n_{1}+n_{2}=10$) and 3 layers, which means that features with lagged as $t-1$, $t-2$, and $t-3$ will be used to predict $RV_{t}$. As mentioned before, we implement two quantum reservoir computing models, QR1 and QR2. QR1 does not utilize the ensemble reservoir approach, while QR2 incorporates an additional reservoir to enhance computational capacity (see Fig. 3). The quantum reservoir outputs, corresponding to evolved quantum states driven by lagged input encodings, are used as nonlinear regressors in a regularized linear model estimated via ridge regression. In both quantum reservoir models, the weight matrix $W_{out}$ is estimated using ridge regression, as given in Eq. (41).
IV.3 Feature selection for quantum reservoir computing

### v2

@@ II.2 Machine Learning Preliminaries
II.2.1 Feedforward Neural Networks
@@
II.2.2 Long Short-Term Memory Neural Networks
@@
where $W_{f}$ is the weight matrix for the forget gate, $\bm{b}_{f}$ is the bias term, and $\bm{f}_{t}$ is a vector whose elements take values between 0 and 11. The input gate determines which parts of the current input $\bm{x}_{t}$ should be added to the cell state $\bm{c}_{t}$. It has two components, a sigmoid layer to decide which values to update, and a tanh layer to create new candidate values to potentially add to the cell state $\bm{c}_{t}$, which formula are defined as
@@
where $\bm{i}_{t}$ is the input gate output, a vector of values between 0 and 11, and $\tilde{\bm{c}}_{t}$ is the candidate cell state. The information from the forget gate and the input gate are used to update the cell state $\bm{c}_{t}$, as
@@
In general, the output $\hat{\bm{y}}_{t}$ is a function of $h_{t}$ as $\hat{\bm{y}}_{t}=g(h_{t})$, where $g(\cdot)$ depends on the specific problem. Training is carried out by optimizing a loss function, of the form of Eq. (5), which aims to minimize the difference between $\hat{\bm{y}}_{t}$ and the real output $\bm{y}_{t}$. During the training procedure, the weight matrices $\{W_{f},W_{i},W_{c},W_{o}\}$ and bias vectors $\{\bm{b}_{f},\bm{b}_{i},\bm{b}_{c},\bm{b}_{o}\}$ are all trained and updated iteratively, see Fig. 1(b). The LSTM architecture uses gating mechanisms, see Fig.1(b), to regulate information flow and gradient propagation [72], effectively mitigating vanishing and exploding gradients common in standard recurrent neural networks and enabling the retention of long-term information. These characteristics make LSTM widely used in fields such as natural language processing and time-series prediction.
II.2.3 Classical Reservoir Computing
@@ III Quantum Reservoir Computing for Realized Volatility Forecasting
Figure 2: Quantum reservoir computing schematic. The task is to forecast $\widehat{RV}_{t}$ based on the past information of features $\bm{x}$ at discrete time from t3t{-}3 to t1t{-}1. The Quantum Reservoir consists of two subsystems, the input qubits $\rho_{I}$ for inputting features and the hidden qubits $\rho_{h}$ for hidden state, composed of $n_{1}$ and $n_{2}$ qubits respectively. In the first, the feature variables $\bm{x}_{t-3}$ are encoded into the input qubits via single-qubit gates $R_{Y}(\bm{x}_{t-3})$ applied on $|0\rangle^{\otimes n_{1}}$. Therefore the input density matrix is given as $\rho_{I}(\bm{x}_{t-3})$. The hidden state at first is initialized as $\rho_{h}(t{-}3)=|0\rangle\langle 0|^{\otimes n_{2}}$. Thus the quantum reservoir state is given by $\rho_{I}(\bm{x}_{t-3})\otimes\rho_{h}(t{-}3)$. The state evolves under the action of the Hamiltonian $H$ for time $\tau$, described by $e^{-iH\tau}$. To update the hidden state for the next step, the input $n_{1}$ qubits are discarded such that the new hidden state is given by a partial trace as $\rho_{h}(t{-}2)$. This process continues until all features from t3t{-}3 to t1t{-}1 have been encoded within the quantum reservoir. After processing all features, measurements on Pauli-Z basis of all qubits are performed to read out useful information from the quantum reservoir, yielding the expectation value $\langle Z_{j}\rangle_{\tau}$ for the $j$-th qubit. These expectation values collectively form the readout vector $\bm{m}_{t}^{(1)}=[\langle Z_{1}\rangle_{\tau},\cdots,\langle Z_{n_{1}+n_{2}}\rangle_{\tau}]$. A linear regression model is used to predict $\widehat{RV}_{t}$ based on this readout vector as $\widehat{RV}_{t}=W_{out}\bm{m}_{t}^{(1)}$, where $W_{out}$ is a weight matrix trained using the ridge regression method.
@@
where $X_{i}$ and $Z_{i}$ are the Pauli operators acting on $i$-th qubits (see Table 4 for the definition of Pauli matrices), $v$ is the strength of the magnetic field which is used as the unit of energy and thus set to be $v=1$ and finally $J_{ij}$ are exchange couplings between qubits $i$ and $j$ which are randomly sampled from $J_{ij}/v\in[0,1]$. After being randomly initialized, the Hamiltonian $H$ is fixed throughout the process, similar to a classical reservoir.
@@
The goal in this paper is to predict $RV_{t}$ by using past $k$ steps features, i.e., the time series $\bm{x}_{t-k},\cdots,\bm{x}_{t-1}\}$. For convenience of notation, we denote $\bm{x}_{t-k}=[x_{t-k,1},x_{t-k,2},\cdots,x_{t-k,n_{1}}]^{T}$, where each element $x_{t-k,i}$ represents the $i$-th economic feature at time step tkt{-}k. For example the input vector $\bm{x}_{t-k}$ may contain three input data at time tkt-k such as realized volatility $RV_{t-k}$, Dividend Yield ($DP_{t-k}$) ratio and Earning-Price ($EP_{t-k}$) ratio (note that all external features used in this paper are defined in Table. 1). In this situation, $\bm{x}_{t-k}=[RV_{t-k},DP_{t-k},EP_{t-k}]^{T}$ contains three input features. In this work, we limit k=3k{=}3, that is, we consider the memory depth for learning to be only three steps. Although generalizing to other memory depths is straightforward and it may appear that considering large memories may yield better results, one must also consider the fact that larger memory depths require deeper quantum circuits. Given the finite coherence times available in near-term quantum devices, this trade-off quickly becomes considerable. As we shall see, our choice of k=3k{=}3 already yields a satisfactory predictive accuracy. Note that all features are scaled between $\pi,+\pi]$, fulfilling the requirement of the $R_{Y}$ gate. The input vectors $\{\bm{x}_{t-3},\bm{x}_{t-2},\bm{x}_{t-1}\}$ are fed into the reservoir model sequentially. The reservoir performs an iterative loop consisting of three steps:
@@
$\text{where }\quad R_{Y}(\bm{x}_{t-3})$ $=$ $\bigotimes_{j=1}^{n_{1}}e^{-i\frac{x_{t-3,j}}{2}Y_{j}}.$ (15)
Rotation encoding is a common choice for near-term quantum hardware because it is easy to implement and embeds classical features into quantum states via parameterized single-qubit rotations, thereby effectively leveraging the continuous degrees of freedom of a qubit state. Therefore, the input density matrix is given by $\rho_{I}(\bm{x}_{t-3})=|\psi_{I}(\bm{x}_{t-3})\rangle\langle\psi_{I}(\bm{x}_{t-3})|$. The hidden $n_{2}$ qubits (the total reservoir thus consists of $n_{1}+n_{2}$ qubits) at this stage are initialized as
@@
Thus, the collective input state of the quantum reservoir state is given by $\rho_{I}(\bm{x}_{t-3})\otimes\rho_{h}(t{-}3)$. Then the whole reservoir evolves under the action of the Hamiltonian $H$ for a specific time $\tau$, described by $U=e^{-iH\tau}$. The evolution scrambles the input data $\bm{x}_{t-3}$ across the entire reservoir. Note that since the reservoir is a fully connected graph we set $\tau=1/v$ which is enough to guarantee that the information is scrambled across the whole system. Following this, while the original $n_{1}$ qubits are discarded, the quantum state of the hidden $n_{2}$ qubits carry this information forward to the next step. The quantum state of the hidden qubits after discarding the input qubits now becomes
@@
$\text{where }\quad R_{Y}(\bm{x}_{t-2})$ $=$ $\bigotimes_{j=1}^{n_{1}}e^{-i\frac{x_{t-2,j}}{2}Y_{j}}.$ (19)
@@
$\text{where }\quad R_{Y}(\bm{x}_{t-1})$ $=$ $\bigotimes_{j=1}^{n_{1}}e^{-i\frac{x_{t-1,j}}{2}Y_{j}}.$ (22)
@@
where $M_{1}=[\cdots,\bm{m}_{t}^{(1)},\cdots,\bm{m}_{T}^{(1)}]$, II is the identity matrix and $\delta{=}10^{-8}$ is a small number which is used to guarantee that the matrix is non-singular. We call this model as Quantum Reservoir 1 (QR1), since here we only use one quantum reservoir. see Appendix B for more information.
@@
Furthermore, to obtain richer information from the quantum reservoir, we adopt the ensemble reservoir approach [40]. In this approach, at any given step $t$ apart from the measurement outcomes $\bm{m}^{(1)}_{t}$, one can use another reservoir which is almost identical to the previous setup except that the evolution at the step III is run for a time duration of $\tau/2$, instead of $\tau$, see Fig. 3.
The ensemble of these two reservoir are combined to make a larger vector $\bm{m}_{t}^{(2)}=[\langle Z_{1}\rangle_{\tau},\ldots,\langle Z_{n_{1}+n_{2}}\rangle_{\tau},\langle Z_{1}\rangle_{\tau/2},\ldots,\langle Z_{n_{1}+n_{2}}\rangle_{\tau/2}]^{T}$, where $\langle Z_{j}\rangle_{\tau/2}$ represents the measurement of the Pauli operator $Z_{j}$ in the second reservoir setup in which the last evolution is run for time $\tau/2$, as schematically shown in Fig. 3. We call this model Quantum Reservoir 2 (QR2) which is expected to be more precise than the QR1 as it uses twice as much resources. It is worth noting that, in the simulation stage, we directly compute the measurement outcomes of the quantum system. In an actual experiment, however, one must perform many repeated measurements (shots) and use their average as the output. Recent studies [110] have shown that, as the system dimension increases, QRC may suffer from an exponential concentration of measurement results toward a value that is independent of the input. Concretely, the variance of measurement outcomes across different input variables can decay exponentially to zero. As a result, an exponential number of shots would be required to reliably distinguish the outputs corresponding to different inputs, which undermines the potential advantage of QRC. In our approach, by contrast, the QRC system is kept at a fixed size (10 qubits), so this issue does not arise (see Appendix C for more information). The measurement outputs corresponding to different inputs still exhibit sufficiently large variance, indicating that our scheme does not require an excessive number of measurements to obtain accurate output estimates.
@@ IV.2 Training and Benchmarking
We employ our quantum reservoir models QR1 and QR2 to predict realized volatility. To assess their performance, we conduct a comprehensive benchmarking study against established classical models. In particular, we compare quantum reservoir computing models QR1 and QR2 with the classical linear models AR1 (which is AR with memory of one previous step), AR3 (which is AR with memory of three previous steps), ARMAX, HAR, HARX as well as the non-linear machine learning based methods LSTM, LSTMX, RC and RCX, see the relevant parts of section II for a short introduction to these models. For all models, including classical models and quantum reservoir computing, we employ a rolling-window approach to estimate and train the models, optimize their parameters, and perform one-step-ahead and five-steps-ahead out-of-sample predictions. Rolling re-estimation at each monthly step is consistent with standard out-of-sample evaluation practices in the volatility forecasting literature, ensuring alignment with recent information and accommodating structural change [117, 118, 58, 119]. This design promotes comparability across models by isolating learning effects from differences in estimation windows and is particularly important for monthly realized volatility, which reflects aggregated market activity and is influenced by slowly evolving macroeconomic regimes. Without regular updating, parameter estimates may become obsolete in the presence of structural breaks. Our approach, therefore, maintains regime adaptiveness while avoiding excessive sensitivity to high-frequency noise. While our study uses monthly updates to maximize predictive accuracy for benchmarking, a lower re-estimation frequency could be adopted in industrial settings to enhance stability and ease of debugging.
@@
The initial training window spans February 1950 to June 1997 (approximately 571 months). After training and optimizing these models on the initial window, we generate a forecast for the next out-of-sample month (August 1997). Subsequently, the rolling window advances by one month, using data from March 1950 to August 1997 to predict September 1997. This process continues iteratively, rolling through all 245 out-of-sample observations, spanning from August 1997 to December 2017. During this process, all models are re-estimated at each step to ensure optimal performance for the new rolling window.
@@
Among the classical linear models that we use, AR1, AR3, ARMAX, HAR, HARX are estimated using ordinary least squares or maximum likelihood methods, depending on the structure of the model. In addition, the machine learning based LSTMX model is implemented with two layers of LSTM cells. The hidden state size is set to 60 for the LSTM model and 50 for the LSTMX variant. After that, a linear regression is used to transform the output of LSTM(X) into the prediction of $\widehat{RV}_{t}$. The model parameters are optimized using the ADAM optimizer, with a learning rate of $0.001$, the batch size of $64$, and $100$ epochs. For the classical reservoir computing models (RC and RCX), the reservoir consists of 50 hidden neurons for RC and 20 hidden neurons for RCX. The leak rate, which controls the speed at which the state of the reservoir evolves, is set at 0.6. The spectral radius, which influences the stability and non-linearity of the reservoir, is fixed at 0.9. The input scaling, which is a coefficient applied on $W_{in}$, is set at $0.1$. The readout matrix $W_{out}$, responsible for mapping the reservoir states to the output, is estimated using ridge regression to mitigate overfitting and improve generalization. All machine learning hyperparameters, including architecture choices and training settings, are selected using time-series-aware cross-validation procedures within the training window to ensure generalization and robustness. (see Appendix D for more informaiotn)
@@
For quantum reservoir computing, taking into account the current limitations of quantum devices, we used a quantum reservoir with 10 qubits (comprising both input and hidden qubits such that $n_{1}+n_{2}=10$) and 3 layers, which means that features with lagged as t1t-1, t2t-2, and t3t-3 will be used to predict $RV_{t}$. As mentioned before, we implement two quantum reservoir computing models, QR1 and QR2. QR1 does not utilize the ensemble reservoir approach, while QR2 incorporates an additional reservoir to enhance computational capacity (see Fig. 3). The quantum reservoir outputs, corresponding to evolved quantum states driven by lagged input encodings, are used as nonlinear regressors in a regularized linear model estimated via ridge regression. In both quantum reservoir models, the weight matrix $W_{out}$ is estimated using ridge regression, as given in Eq. (26). Implementation-wise, we consider 100 different quantum reservoir instances and train and evaluate each of them separately. We then select the best-performing quantum reservoir and report the results. Similarly for LSTM, we test multiple hyperparameter configurations and report the best-performing results for presentation. Notice that we are reporting the best-performing results for both LSTM and QRC, hence QRC does not specifically gain any unfair advantage.
IV.3 Closed-Loop Prediction of Multi-Step Future Outcomes
So far, we have focused on the prediction of outcomes which are only one step ahead of the input. In other words, all models are trained under a one-step open-loop setting: the model inputs are ground-truth values, and performance is evaluated against the ground truth. For example,
$\widehat{RV_{t}}=f\!\left(RV_{t-3},Xs_{t-3},\,RV_{t-2},Xs_{t-2},\,RV_{t-1},Xs_{t-1}\right),$
where XsXs denotes other (exogenous) variables (see Table 1). If a model name does not include the X suffix, then XsXs is empty (i.e., no exogenous variables are used). This is called open-loop strategy in which historic data is used for predicting one step ahead. However, one might be interested to predict the outcomes well ahead of the current data. For instance, by accessing the historic data one might be interested to know the outcome of $S$ steps ahead. This can be done through a closed-loop procedure which is explained below. To accomplish this, we rely on a conventional trained open-loop model which predict $S=1$ step ahead. For predicting $S>1$ steps in the future, we repeatedly call the model and feed its previous predictions back to the model as part of the next input. Specifically:
  • For $S=1$,
    $\widehat{RV_{t}}=f\!\left(RV_{t-3},Xs_{t-3},\,RV_{t-2},Xs_{t-2},\,RV_{t-1},Xs_{t-1}\right).$
  • For $S=2$,
    $\widehat{RV_{t+1}}=f\!\left(RV_{t-2},Xs_{t-2},\,RV_{t-1},Xs_{t-1},\,\widehat{RV_{t}},Xs_{t}\right).$
    At this step, $\widehat{RV_{t}}$ is taken from the previous step’s output. Since our models predict only RVRV (and not the other variables), $Xs_{t}$ is still provided as ground truth. However, if such data is not available then a new predictor has to be trained targeting $Xs_{t}$ and use its predicted value (i.e. $\widehat{Xs}_{t}$) as the input of the closed-loop model. Here, for simplicity we assume that the ground truth of $Xs_{t}$ is available.
Noted that when $S=1$, the closed-loop setting degenerates to the open-loop setting. Using the closed-loop strategy allows us to assess a model’s ability to forecast into the future, which is crucial in many forecasting scenarios. We consider $S=1$ and $S=5$ to evaluate both short-term and longer-term predictive performance. It is worth emphasizing that by increasing $S$ the accuracy is expected to go down as the inputs come from prediction rather than the actual ground truth data. In other words, at every steps, a small error can propagate to the next steps affecting the accuracy of the model for predicting distant future outcomes.
IV.4 Feature Selection for Quantum Reservoir Computing